(a) A sphere has a surface uniformly charged with 1.00 C. At what distance from its center is the potential 5.00 MV? (b) What does your answer imply about the practical aspect of isolating such a large charge? Show the process.


1
Expert's answer
2022-01-19T06:01:04-0500

(a) The distance can be found from the given data:


"V=\dfrac{Q}{4\pi\epsilon_0r}→r=\dfrac{Q}{4\pi\epsilon_0V}=1798\text{ m}."


(b) Such a large charge requires a lot of space to be isolated, which can be hardly implemented practically. A Faraday cage of such a large size is very hard to build practically.